OpenStudy (sonyalee77):

Which of the following are point-slope equations of the line going through (6, 7) and (2, -1)?

OpenStudy (sonyalee77):

@eliassaab @wolfe8

OpenStudy (anonymous):

Where are your choices?

OpenStudy (sonyalee77):

slope 9/5 , -9/5,5/9,or-5/9

OpenStudy (anonymous):

(6, 7) and (2, -1)? \[ slpoe=\frac { 7-(-1)}{6-2}=\frac {8}{4}=2 \]

OpenStudy (anonymous):

None of your choices are correct.

OpenStudy (sonyalee77):

found out the answer thanx